BRACEFACEJAH:

x/2x+8=3/x+2...SOLVE FOR X

Shadow:

\[\frac{ x }{ 2x + 8 } = \frac{ 3 }{ x + 2 }\]

Shadow:

The fractions make this a bit awkward to solve, so what do you think would be your first step?

Shadow:

Well you could start by removing the denominators so that you don't have to divide.

Shadow:

Cross multiplication is a pretty simple way to do this. Basically do:\[x \times (x + 2) = ?\] and \[3 \times (2x + 8)\]

Shadow:

@bracefacejah Does that make sense from there?

BRACEFACEJAH:

yes thnks

BRACEFACEJAH:

do i add after that?

Shadow:

After this you should get: \[x^2 + 2x = 6x + 24\]

Shadow:

Looking at this you know you're going to have to use the quadratic formula to solve for this, because of the x^2, so basically you want to create ax + by + c = 0, so get all the terms to the left side.

Shadow:

Subtract 6x from both sides, then 24 from both sides, you get \[x^2 - 4x - 24 = 0\]

Shadow:

If you don't know the quadratic formula, this is it: \[x = \frac{ -b \pm \sqrt (b^2 - 4ac) }{ 2a }\] Where a quadratic equation looks like, ax^2 + by + c = 0 So from x^2 - 4x - 24 we can grab the coefficients, a = 1, b = -4 c = -24 Then you just input these into the formula and solve for x.

Shadow:

Does that make sense so far?

BRACEFACEJAH:

yes

Shadow:

\[x = \frac{ -(-4) \pm \sqrt(-4)^2 - 4(1)(-24) }{ 2(1) }\]

BRACEFACEJAH:

i got my answer... -5

Shadow:

Are you sure?

BRACEFACEJAH:

yes

Shadow:

You can get very different answers based off of how you look at the fractions.

BRACEFACEJAH:

i checked my answer and its right

Shadow:

Like is it (x/2x) +8=(3/x)+2 or x/2x+8=3/x+2

BRACEFACEJAH:

no it was -5

Shadow:

Frankly, I don't see how you get 5 from this. I double checked it with Wolfram. Are you sure you gave me the right problem lol?

BRACEFACEJAH:

yes i gave the right question and I put it into the system I'm getting questions on and it says its right

Shadow:

Oh okay. It may just be a lesson glitch or something. If you input x/(2x+8)=3/(x+2) into Wolfram, you definitely don't get -5